OpenStudy (anonymous):

4 sin theta + 3 root 3 = root 3

hero (hero):

You are given\[4 \sin(\theta) + 3\sqrt{3} = \sqrt{3}\] and need to solve for \(\theta\) correct?

OpenStudy (anonymous):

yes

hero (hero):

First, subtract \(3\sqrt{3}\) from both sides.

OpenStudy (anonymous):

did that

hero (hero):

@MRVerdi, show what you got afterwards.

OpenStudy (anonymous):

4 sin theta = 3?

hero (hero):

Let me ask you something. Does x - 3x = 3?

OpenStudy (anonymous):

no x=-3/2

hero (hero):

You are very confused

OpenStudy (anonymous):

....

hero (hero):

What is 1 - 3 ?

OpenStudy (anonymous):

-2

hero (hero):

Okay so... x - 3x = (1 - 3)x = -2x

OpenStudy (anonymous):

if you solving for x then you have to divide by -2

hero (hero):

Similarly \(\sqrt{3} - 3\sqrt{3} = (1 - 3)\sqrt{3} = -2\sqrt{3}\)

hero (hero):

There is something about this that you are confused about. You're not solving for x. You are simply subtracting.

OpenStudy (anonymous):

sorry i misunderstood what you were asking

OpenStudy (anonymous):

then how do you divide the -2 root 3 by 4?

hero (hero):

You should end up with \[\sin(\theta) = -\frac{2\sqrt{3}}{4}\] Now take the inverse sine of both sides to isolate \(\theta\)

OpenStudy (anonymous):

\[4 \sin \theta + 3 \sqrt 3 = \sqrt 3 \implies 4 \sin \theta = -2 \sqrt 3 \implies \sin \theta =- \frac{\sqrt3}{2}\]\[\implies \theta= - \frac {\pi}{3} + 2k \pi ~~or~~ \theta = - \frac {2\pi}{3} + 2k \pi\]

OpenStudy (anonymous):

Depending on your text, the first term of the two answers might have to be positive as a matter of convention. That would mean the answers would look like \[\theta = \frac{4 \pi}{3} + 2k \pi ~~or~~ \theta = \frac{5 \pi}{3} + 2k \pi\] The main thing about this problem is to work it to a familiar form. Then you HAVE to recognize the special values of the functions, and know where on the unit circle they occur.

hero (hero):

@AnimalAin, if you noticed, I was trying to help the user solve it step by step.

OpenStudy (anonymous):

The plus 2k pi takes into account that every time you come past that point on the unit circle, you will get a correct answer. Note that k is any integer.

OpenStudy (anonymous):

Sorry if I interrupted.